There are 89 students in senior year out of whom all but 3 study latin, spaniah or french. If 25 take latin, 32 spanish and 37 french and no student takes all three languages, how many students take more than one language?
A)0
B)3
C)5
D)8
E)10

Answer by stanbon(75887)

There are 89 students in senior year out of whom all but 3 study latin, spaniah or french. If 25 take latin, 32 spanish and 37 french and no student takes all three languages, how many students take more than one language?
----
Draw the Venn Diagram of 3 intersecting circles.
Label the circles latin, span, and french.
----
The universe has 89 elements.
3 elements are not in any of the circles, leaving 86 in the circles.
----
latin has 25 elements
span has 32 elements
french has 37
The intersection of the 3 circles has 0 elements
----
(25+32+37)-86 = 94-86 = 8 students take more than one language.
============
Cheers,
Stan H.
===============
A)0
B)3
C)5
D)8
E)10